A company produces a product at a cost of 6 rials per unit. The fixed costs are 2000 rials.
The product sells at 8 rials per unit.
(a) Find the total cost YT C.
(b) Find the total revenue YT R.
(c) Find how many units must be produced and sold to make a profit of 12000 rials.

Answer by Theo(13342)

c = 2000 + 6x.

r = 8x.

x represents the number of units produced and sold.
c represents the total cost of production.
r represents the revenue.

p = r - c

p represents the total profit.

replace r with 8x and c with 2000 + 6x and the profit formula becomes:

p = 8x - (2000 + 6x)

simplify to get p = 8x - 6x - 2000

combine like terms to get p = 2x - 2000

replace p with 12000 to get 12000 = 2x - 2000

add 2000 to both sides of the equation to get 14000 = 2x

solve for x to get x = 14000 / 2 = 7000.

he would have to produce and sell 7000 units to make a profit of 12000 rials.

when x = 7000, .....

c = 2000 + 6 * 7000 = 44000 rials.

r = 8 * 7000 = 56000 rials.

p = r - c = 56000 - 44000 = 12000 rials.

your solution is that 7000 units must be produced and sold to make a profit of 12000 rials.